OpenStudy (anonymous):

Please!!!! Help Ada is showing her work in simplifying (5.9 – 1.8) – 6.2 + 4.1. In which step did Ada make an error? Step 1: (5.9 – 1.8) – 6.2 + 4.1 Step 2: 5.9 + (–1.8 – 6.2) + 4.1 (associative property) Step 3: 5.9 – 8 + 4.1 Step 4: 5.9 + 4.1 – 8 (associative property) Step 5: 10 – 8 = 2 A). Step 2; she wrote associative instead of commutative B). Step 2; she wrote associative instead of distributive C). Step 4; she wrote associative instead of commutative D). Step 4; she wrote associative instead of distributive

OpenStudy (ninjaslice):

B

OpenStudy (anonymous):

i dont think its B? how do you know??

OpenStudy (anonymous):

ello love :D

OpenStudy (anonymous):

Hey @atreyu6s3x6 <3

OpenStudy (anonymous):

need some help?

OpenStudy (anonymous):

yes I do :(

OpenStudy (anonymous):

im am not positive that the answer is b

OpenStudy (anonymous):

i was heading for c? but i want to be positive

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

sorry love, slow internet haha but lets do this step by step :)

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Associative Property: The associative property states that you can add or multiply regardless of how the numbers are grouped. By 'grouped' we mean 'how you use parenthesis'. In other words, if you are adding or multiplying it does not matter where you put the parenthesis. Step 4 doesn't deal with associative property, it deals with the commutative property. Commutative Property: The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2. @atreyu6s3x6

OpenStudy (anonymous):

<3 :)

OpenStudy (anonymous):

seems correct haha sorry my interwebs are slow as hell :P

OpenStudy (anonymous):

Haha But thanks anyways <3 :)

OpenStudy (anonymous):

no problem love, tagg me anytime :)